# Active Fine-Tuning of Generalist Policies

### v1

@@ 1 Introduction
The availability of large pre-trained models has transformed entire areas of machine learning, from computer vision (Krizhevsky et al., 2012; He et al., 2016; Dosovitskiy et al., 2021; Radford et al., 2021), to natural language processing (Radford et al., 2019; Brown et al., 2020) and generative modeling in general (Ho et al., 2020; Esser et al., 2024). This paradigm has started to extend to robotics and control (Collaboration, 2023; Ma et al., 2024), in particular for systems for which demonstrations are readily available (Octo Model Team et al., 2024), or can be easily collected (Zhao et al., 2023). Even when demonstrations are not easily obtained, scaling laws in reinforcement learning (Ceron et al., 2024b; a; Nauman et al., 2024) suggest the possibility of leveraging large pre-trained policies. These “generalist” policies have decent performance on many tasks, and can be fine-tuned on particular set of tasks while leveraging their previously learned representations and skills. We investigate whether representations of such policies can be used to significantly bootstrap learning progress.
As a motivating example, consider a household robot that is delivered with a pre-trained “generalist” policy, and deployed in slightly different conditions than those observed in its training data. While the robot may achieve some tasks in a zero-shot fashion (e.g. simple pick-and-place), other tasks might necessitate further fine-tuning (e.g. cooking an omelette). The robot should be able to interactively request demonstrations to compensate for its shortcomings. We seek to answer which demonstrations should be requested from the user to achieve the best performance, as quickly as possible.
@@
To address this problem, we propose AMF (Active Multi-task Fine-tuning), which selects maximally informative demonstrations. AMF parallels recent work on active supervised fine-tuning of neural networks (Hübotter et al., 2024). To this end, AMF relies on estimates of the demonstrations’ information gain about the expert policy. We prove that in sufficiently regular Markov decision processes, AMF converges to the expert policy. We then focus on practical scenarios where policies are represented as neural networks. We show that, despite additional challenges, AMF can effectively guide active task selection in such settings, leading to better policies after fewer demonstrations. Our contributions are:
@@ 2 Related Work
Learning-based control and active data selection are both well-established research directions. This section discusses some of the most topical works in either direction, and clarifies the novelty and placement of this work with respect to them.
Behavioral Cloning
@@
Numerous imitation learning approaches have been developed with the goal of distilling knowledge from high-quality demonstrations to a control policy (Osa et al., 2018). Within this family of techniques, behavioral cloning (BC, Bain & Sammut, 1995; Ross & Bagnell, 2010) aims to maximize policy performance by minimizing the distance of its actions to demonstrated actions, simply through supervised learning. While BC may suffer from accumulating errors (Ross et al., 2011), its empirical effectiveness has seen increasing support when high-quality demonstrations are readily available (Kumar et al., 2022). Next to recent empirical successes (Chi et al., 2023), formal analysis has also advanced (Spencer et al., 2021; Block et al., 2024a; Belkhale et al., 2024; Foster et al., 2024), and established provable performance guarantees for BC policies (Xu et al., 2020; Maran et al., 2023; Block et al., 2024b).
Multi-task and Generalist Policies
@@
Traditionally, behavioral cloning has mostly been deployed in a single-task setting. Multi-task learning in sequential decision-making has largely been investigated in the context of reinforcement learning (Teh et al., 2017; Sodhani et al., 2021; Yu et al., 2021; Sun et al., 2022; Cho et al., 2022; Hendawy et al., 2023). Moreover, the recent rise of multi-task generative models (Brown et al., 2020) has been mirrored by exploration of multi-task, or generalist policies, often trained via imitation learning (Reed et al., 2022; Bousmalis et al., 2024; Collaboration, 2023). These recent works mostly build upon algorithms developed for the single-task case, and simply integrate task-conditioning as part of the state. While several works hand-select parts of large, open-source robotics datasets for pre-training (Octo Model Team et al., 2024), active data selection for multi-task fine-tuning has not been addressed. Prior work on meta-learning has studied how one can explicitly meta-learn the ability to adapt to task demonstrations (Finn et al., 2017). We find this capability to emerge even from models that are not explicitly trained in this way, and focus on which demonstrations to obtain.
Data Selection
@@
The idea of directing a sampling process to gather information has been central to machine learning research and studied extensively in experimental design (Chaloner & Verdinelli, 1995) and active learning (Settles, 2009). Most work on active data selection summarizes data without focusing on a particular task (e.g., Sener & Savarese, 2017; Ash et al., 2020b; Holzmüller et al., 2023; Lightman et al., 2023), which has been predominantly applied to pre-training. Recently, adapting models after pre-training and during deployment has gained interest. Several works, mostly in computer vision, focus on unsupervised fine-tuning on a test instance (Jain & Learned-Miller, 2011; Krause et al., 2018; Sun et al., 2020; Wang et al., 2021b; Chen et al., 2022). We focus instead on supervised fine-tuning of learning-based controllers in dynamical systems. This necessitates automatic data selection, for which practical methods currently rely on uniform sampling or externally provided heuristics. Our approach extends work on task-directed data selection (Kothawade et al., 2020; Wang et al., 2021a; Kothawade et al., 2022; Smith et al., 2023), which has recently been applied to the supervised fine-tuning of large-scale neural networks in vision (Hübotter et al., 2024) and language (Xia et al., 2024).
@@ 3.1 Multi-Task Reinforcement Learning
$\bm{\tau}_{\bm{\pi}}\Big{(}\big{(}s_{0},a_{0},s_{1},a_{1},\dots\big{)}\mid c\Big{)}=\mu_{0}(s_{0})\prod_{t=0}^{\infty}\bm{\pi}(a_{t}\mid s_{t},c)\cdot P(s_{t+1}\mid s_{t},a_{t}).$
@@
Reinforcement learning algorithms traditionally aim directly at maximizing $J_{\mu_{c}}^{\bm{\pi}}$, which is notoriously challenging. In the scope of this work, we instead consider an imitation learning setting, in which expert demonstrations from an optimal policy $\pi^{\star}$ are provided. In particular, we focus on behavioral cloning algorithms, which reduce the control problem to a supervised learning problem. Given a set of $N$ task-conditioned, $H$-length trajectories $\hat{\tau}_{1:N}=(s_{0}^{i},a_{0}^{i},\dots,s_{H-1}^{i},a_{H-1}^{i})_{i=1}^{N}$ with task labels $c_{1:N}$, behavioral cloning proposes a proxy objective for the policy $\pi$: an empirical estimate of the log-likelihood under the data distribution: $J^{\bm{\pi}}_{\text{proxy}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{H-1}\log\bm{\pi}(a_{t}^{i}\mid s_{t}^{i},c_{i}).$ If trajectories $\tau_{1:N}$ are obtained from the optimal policy, cover the support of the desired task distribution $\mu_{c}$, and the searched policy class is sufficiently rich, the maximizer of $J^{\bm{\pi}}_{\text{proxy}}$ will also maximize $J^{\bm{\pi}}_{\mu_{c}}$ as $N$ and $H$ increase. However, in general, there is a clear mismatch between $J^{\bm{\pi}}_{\text{proxy}}$ and $J^{\bm{\pi}}_{\mu_{c}}$ (Xu et al., 2020; Maran et al., 2023). Nonetheless, the optimization of $J^{\bm{\pi}}_{\text{proxy}}$ is a relatively straightforward supervised learning problem, while the full RL problem raises several convergence issues, particularly in the offline setting (Levine et al., 2020). Thus, we use $J^{\bm{\pi}}_{\mu_{c}}$ only for evaluation, and carry out optimization through the proxy objective.
@@ 3.2 Active Policy Fine-Tuning
In this work, we consider an active fine-tuning scheme for multi-task policies. The goal is to fine-tune a pre-trained policy to perform well on a desired task distribution $\mu_{c}$ using as few expert demonstrations as possible. The agent is allowed $N$ sequential queries for demonstrations according to the fine-tuning budget. The $n$-th query should consist of a task $c_{n}\in\mathcal{C}$. Once the agent selects a task, feedback is received from the optimal policy $\pi^{\star}:\mathcal{S}\times\mathcal{C}\to\mathcal{A}$ (i.e., an optimal demonstrator). At each round the agent receives an $H$-step demonstration conditioned on the chosen task $c_{n}$. This can be seen as a single measurement from a stochastic process over trajectories $\bm{\tau}:\mathcal{C}\to\Delta((\mathcal{S}\times\mathcal{A})^{H})$.
Each observed trajectory up to round $n$ is stored in a dataset $(c_{1:n},\hat{\tau}_{1:n})$, which can be used to fine-tune the policy, and condition the agent’s query at step $n+1$. The process is repeated for $N$ rounds, with the goal of producing a fine-tuned policy that maximizes the expected returns for the desired task distribution $\mu_{c}$.
Modeling assumptions
We take a Bayesian perspective on active multi-task fine-tuning, by assuming a Bayesian model $\bm{\pi}$ over policies. We assume that demonstrations follow a noisy expert: $\bm{\tilde{\pi}}(s,c)=\pi^{\star}(s,c)+\bm{\epsilon}(s,c)$ where $\bm{\epsilon}(s,c)$ is independent noise. We remark, however, that AMF can also be understood from a non-Bayesian perspective as selecting tasks that most quickly minimize the size of frequentist confidence sets around the optimal policy.
@@ 4 Method
We show in Section 4.1 that, under certain regularity assumptions, the policy learned by AMF converges to the expert policy and matches its performance. These results constitute a first-of-its-kind performance guarantee for active multi-task fine-tuning. The main novelty of this guarantee is the extension of prior work to sequential domains where the visited trajectory $(s_{0},a_{0},s_{1},\dots)\sim\bm{\tau}(c_{n})$ is unknown when selecting the task $c_{n}$ for a demonstration. In Section 4.2, we discuss the design choices that make AMF amenable to optimization in practical settings.
@@ 4.1 Performance Guarantees
Informal Assumption 1.
We make the following assumptions:
@@
Informal Theorem 2 (Performance guarantees for active multi-task BC).
@@
$J^{\pi^{\star}}_{\mu_{c}}-J^{\pi_{n}}_{\mu_{c}}\leq O(\gamma_{(Hn)}n^{-\frac{1}{2}}),$
@@
Intuitively, this theorem proves that the imitator will eventually achieve the demonstrator’s performance in smooth, regular MDPs with sublinear $\gamma_{(Hn)}$ (for a formal definition, we refer to Lemma 2 in the Appendix). We can also prove a more general result under weaker assumptions: as long as the policy is regular, the imitator will reach the noisy expert performance in arbitrary, non-smooth MDPs, albeit only in expectation. A full derivation of this further result can be found in Appendix B.
@@ 4.2 Practical Algorithms
Theorem 2 guarantees that, under regularity assumptions, the adaptive demonstration sampling scheme leads to convergence of the imitator’s performance to the optimal one. However, this criterion involves state occupancies and a conditional entropy term, which are hard to access or estimate in practice. Thus, here we derive a practical objective to be deployed in general settings. We first rephrase the objective from Equation 1 in its entropy form:
$c_{n}$ $=\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1}),$ (2)
@@
Computing the expectation over a policy’s occupancy over states or trajectories is in general intractable in continuous state spaces. Fortunately, a coarse empirical estimate can be obtained as soon as few expert demonstrations become available. The expectation $\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}(\cdot)$ can be estimated through a single sample, which is always available in the form of the trajectories $\hat{\tau}_{1:n-1}$ collected so far, as they have effectively been sampled from $\bm{\tau}(c_{1:n-1})$. However, the remaining two expectations (i.e., $\mathop{\mathbb{E}}_{\tau^{\prime}\sim\bm{\tau}(c^{\prime})}(\cdot)$ and $\mathop{\mathbb{E}}_{(s_{0},\dots)\sim\bm{\tau}(c)}(\cdot)$) involve the distribution over trajectories for an arbitrary task, which might not have been demonstrated yet. However, we observe that, at round $n$, the tasks demonstrated so far induce the empirical distribution $\hat{\mu}_{c}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{c_{i}}(\cdot)$, while the trajectories collected similarly induce $\bm{\hat{\tau}}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{\hat{\tau}_{i}}(\cdot)$, where $\delta$ indicates the Dirac delta distribution. We can show that expectations over the trajectory distribution for an arbitrary task $c\in\mathcal{C}$ can be estimated through importance sampling (i.e., by sampling trajectories from $\bm{\hat{\tau}}(\cdot)$ instead of $\bm{\tau}(\cdot|c)$): $\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(\cdot|c)}f(\tau)=\mathop{\mathbb{E}}_{\tau\sim\bm{\hat{\tau}}(\cdot)}\frac{\bm{\tau}(\tau|c)}{\bm{\hat{\tau}}(\tau)}f(\tau)$. The importance weights can then be estimated as
@@
where $\tau=(s_{0},a_{0},\dots)$ and $\bm{\tilde{\pi}}$ can be approximated with the current estimate of $\bm{\pi}$. Intuitively, the likelihood ratio of a trajectory under two different tasks only depends on the likelihood of actions under the policy, and thus does not require knowledge of the MDP. As the estimate may be inaccurate for small numbers of samples, in practice the algorithm can invest the first few rounds to query a single demonstration for each of the tasks (in case they are countable and few) or to sample the task space uniformly. On the other hand, the high-variance of the estimate can be controlled by practical solutions such as clipping. We present the resulting empirical estimate for Equation 2 in full in Appendix C.
@@
     Update $\pi_{n+1}$ from $\pi_{n}$ with $\mathcal{D}_{n}=\mathcal{D}_{n-1}\cup\{c_{n},\tau_{n}\}$
@@
Thus, when the policy can be modeled as a GP, the only approximation needed concerns occupancy estimation. We refer to this first, practical instantiation as AMF-GP, and present a general algorithmic framework in Algorithm 1. Application of the method to policies parameterized by neural networks will adopt the same scheme. It will also require additional care on three distinct topics: (i) kernel approximations, (ii) batch selection and (iii) forgetting mitigation.

### v3

# Active Fine-Tuning of Multi-Task Policies

@@ Abstract
Imitation Learning, Multi-task Learning, Active Learning, Deep Reinforcement Learning
@@ 1 Introduction
The availability of large pre-trained models has transformed entire areas of machine learning, from computer vision (Krizhevsky et al., 2012; He et al., 2016; Dosovitskiy et al., 2021; Radford et al., 2021), to natural language processing (Radford et al., 2019; Brown et al., 2020) and generative modeling in general (Ho et al., 2020; Esser et al., 2024). This paradigm has started to extend to robotics and control (Collaboration, 2023; Ma et al., 2024), in particular for systems for which demonstrations are readily available (Octo Model Team et al., 2024), or can be easily collected (Zhao et al., 2023). Even when demonstrations are not easily obtained, scaling laws in reinforcement learning (Ceron et al., 2024b, a; Nauman et al., 2024) suggest the possibility of leveraging large pre-trained policies. These “generalist” policies have decent performance on many tasks, and can be fine-tuned on particular set of tasks while leveraging their previously learned representations and skills. We investigate whether representations of pre-trained policies can be used to significantly bootstrap learning progress.
@@
As a motivating example, consider a household robot that is delivered with a pre-trained “generalist” policy, and deployed in different conditions than those observed in its training data. While the robot may achieve some tasks in a zero-shot fashion (e.g. simple pick-and-place), other tasks might necessitate further fine-tuning (e.g. cooking an omelette). The robot should be able to interactively request demonstrations to compensate for its shortcomings. We seek to answer which demonstrations should be requested from the user to achieve the best performance, as quickly as possible.
@@
To address this problem, we propose AMF (Active Multi-task Fine-tuning), which selects maximally informative demonstrations. AMF parallels recent work on active supervised fine-tuning of neural networks (Hübotter et al., 2024). To this end, AMF relies on estimates of the demonstrations’ information gain on the expert policy. We prove that in sufficiently regular Markov decision processes, AMF converges to the expert policy. We then focus on practical scenarios where policies are represented as neural networks. We show that, despite additional challenges, AMF can effectively guide active task selection in such settings, leading to better policies after fewer demonstrations. Our contributions are:
@@
  • We additionally propose a practical approach to alleviate catastrophic forgetting in pre-trained neural policies, which benefits arbitrary data selection strategies.
@@ 2 Related Work
Learning-based control and active data selection are both well-established research directions. This section discusses some topical works in either direction, and clarifies the novelty and placement of this work with respect to them.
@@
Behavioral Cloning  Numerous imitation learning approaches have been developed with the goal of distilling knowledge from high-quality demonstrations to a control policy (Osa et al., 2018). Within this family of techniques, behavioral cloning (BC, Bain & Sammut, 1995; Ross & Bagnell, 2010) aims to maximize policy performance by minimizing the distance of its actions to demonstrated actions, simply through supervised learning. While BC may suffer from accumulating errors (Ross et al., 2011), its empirical effectiveness has seen increasing support when high-quality demonstrations are readily available (Kumar et al., 2022). Next to recent empirical successes (Chi et al., 2023), formal analysis has also advanced (Spencer et al., 2021; Block et al., 2024a; Belkhale et al., 2024; Foster et al., 2024), and established provable performance guarantees for BC policies (Xu et al., 2020; Maran et al., 2023; Block et al., 2024b).
@@
Multi-task and Generalist Policies  Traditionally, behavioral cloning has mostly been deployed in a single-task setting. Multi-task learning in sequential decision-making has largely been investigated in the context of reinforcement learning (Teh et al., 2017; Sodhani et al., 2021; Yu et al., 2021; Sun et al., 2022; Cho et al., 2022; Hendawy et al., 2023). Moreover, the recent rise of multi-task generative models (Brown et al., 2020) has been mirrored by exploration of multi-task, or generalist policies, often trained via imitation learning (Reed et al., 2022; Bousmalis et al., 2024; Collaboration, 2023). These recent works mostly build upon algorithms developed for the single-task case, and simply integrate task-conditioning as part of the state. While several works hand-select parts of large, open-source robotics datasets for pre-training (Octo Model Team et al., 2024), active data selection for multi-task fine-tuning has not been addressed. Prior work on meta-learning has studied how one can explicitly meta-learn the ability to adapt to task demonstrations (Finn et al., 2017). We find this capability to emerge even from models that are not explicitly trained in this way, and focus on which demonstrations to obtain.
@@
Data Selection  The idea of directing a sampling process to gather information has been central to machine learning research and studied extensively in experimental design (Chaloner & Verdinelli, 1995), active learning (Settles, 2009) and reinforcement learning (Mehta et al., 2022). Most work on active data selection summarizes data without focusing on a particular task (e.g., Sener & Savarese, 2017; Ash et al., 2020; Holzmüller et al., 2023; Lightman et al., 2023), which has been predominantly applied to pre-training. Recently, adapting models after pre-training and during deployment has gained interest. Several works, mostly in computer vision, focus on unsupervised fine-tuning on a test instance (Jain & Learned-Miller, 2011; Krause et al., 2018; Sun et al., 2020; Wang et al., 2021b; Chen et al., 2022). We focus instead on supervised fine-tuning of learning-based controllers in dynamical systems. Our approach extends work on task-directed data selection (Kothawade et al., 2020; Wang et al., 2021a; Kothawade et al., 2022; Bickford Smith et al., 2023), which has recently been applied to the supervised fine-tuning of large-scale neural networks in vision (Hübotter et al., 2024) and language (Rotman & Reichart, 2022; Xia et al., 2024; Hübotter et al., 2025).
@@ 3.1 Multi-Task Reinforcement Learning
$\bm{\tau}_{\bm{\pi}}$ $\Big{(}\big{(}s_{0},a_{0},s_{1},a_{1},\dots\big{)}\mid c\Big{)}=$
$\mu_{0}(s_{0})\prod_{t=0}^{\infty}\bm{\pi}(a_{t}\mid s_{t},c)\cdot P(s_{t+1}\mid s_{t},a_{t}).$
@@
Reinforcement learning algorithms traditionally aim directly at maximizing $J_{\mu_{c}}^{\bm{\pi}}$, which is notoriously challenging. In the scope of this work, we instead consider an imitation learning setting, in which expert demonstrations from an optimal policy $\pi^{\star}$ are provided. In particular, we focus on behavioral cloning algorithms, which reduce control to a supervised learning problem. Given a set of $N$ task-conditioned, $H$-length trajectories $\hat{\tau}_{1:N}=(s_{0}^{i},a_{0}^{i},\dots,s_{H-1}^{i},a_{H-1}^{i})_{i=1}^{N}$ with task labels $c_{1:N}$, behavioral cloning proposes a proxy objective for the policy $\bm{\pi}$: an empirical estimate of the log-likelihood under the data distribution: $J^{\bm{\pi}}_{\text{proxy}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{H-1}\log\bm{\pi}(a_{t}^{i}\mid s_{t}^{i},c_{i}).$ If trajectories $\tau_{1:N}$ are obtained from the optimal policy, cover the support of the desired task distribution $\mu_{c}$, and the searched policy class is sufficiently rich, the maximizer of $J^{\bm{\pi}}_{\text{proxy}}$ will also maximize $J^{\bm{\pi}}_{\mu_{c}}$ as $N$ and $H$ increase. However, in general, there is a clear mismatch between $J^{\bm{\pi}}_{\text{proxy}}$ and $J^{\bm{\pi}}_{\mu_{c}}$ (Xu et al., 2020; Maran et al., 2023). Nonetheless, the optimization of $J^{\bm{\pi}}_{\text{proxy}}$ is a relatively straightforward supervised learning problem, while the full RL problem raises several convergence issues, particularly in the offline setting (Levine et al., 2020). Thus, we use $J^{\bm{\pi}}_{\mu_{c}}$ only for evaluation, and carry out optimization through the proxy objective.
@@ 3.2 Active Policy Fine-Tuning
In this work, we consider an active fine-tuning scheme for multi-task policies. The goal is to fine-tune a pre-trained policy to perform well on a desired task distribution $\mu_{c}$ using as few expert demonstrations as possible. The agent is allowed $N$ sequential queries for demonstrations according to the fine-tuning budget. The $n$-th query should consist of a task $c_{n}\in\mathcal{C}$. Once the agent selects a task, feedback is received from the optimal policy $\pi^{\star}:\mathcal{S}\times\mathcal{C}\to\mathcal{A}$ (i.e., an optimal demonstrator). At each round the agent receives an $H$-step demonstration conditioned on the chosen task $c_{n}$. This can be seen as a single measurement from a stochastic process over trajectories $\bm{\tau}:\mathcal{C}\to\Delta((\mathcal{S}\times\mathcal{A})^{H})$. Each observed trajectory up to round $n$ is stored in a dataset $(c_{1:n},\hat{\tau}_{1:n})$, which can be used to fine-tune the policy, and condition the agent’s query at step $n+1$. The process is repeated for $N$ rounds, with the goal of producing a fine-tuned policy that maximizes the expected returns for the desired task distribution $\mu_{c}$. We note that the fine-tuning process does not assume access to the pre-training data distribution. This setting is both realistic, as large pre-training datasets are rarely publicly available, and challenging, as it prohibits any naive data rebalancing strategy.
Modeling assumptions   We take a Bayesian perspective on active multi-task fine-tuning, by assuming a Bayesian model $\bm{\pi}$ over policies. We assume that demonstrations follow a noisy expert: $\bm{\tilde{\pi}}(s,c)=\pi^{\star}(s,c)+\bm{\epsilon}(s,c)$ where $\bm{\epsilon}(s,c)$ is independent noise. We remark, however, that AMF can also be understood from a non-Bayesian perspective as selecting tasks that most quickly minimize the size of frequentist confidence sets around the optimal policy.
@@ 4 Method
$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tau}(c^{\prime})\mid c_{1:n-1},\tau_{1:n-1}),$
$\text{with }c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c),\;\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}).$ (1)
We show in Section 4.1 that, under certain regularity assumptions, the policy learned by AMF converges to the expert policy and matches its performance. These results constitute a first-of-its-kind performance guarantee for active multi-task fine-tuning. The main novelties of this guarantee are the extension of prior work to sequential domains where the visited trajectory $(s_{0},a_{0},s_{1},\dots)\sim\bm{\tau}(c_{n})$ is unknown when selecting the task $c_{n}$ for a demonstration, and the connection to imitator performance. In Section 4.2, we discuss the design choices that make AMF amenable to optimization in practical settings.
@@ 4.1 Performance Guarantees
Informal Assumption 4.1.
We make these assumptions:
@@
Informal Theorem 4.2 (Performance guarantees for active multi-task BC).
@@
$J^{\pi^{\star}}_{\mu_{c}}-J^{\pi_{n}}_{\mu_{c}}\leq O(\gamma_{(Hn)})/\sqrt{n},$
@@
Intuitively, this theorem proves that the imitator will eventually achieve the demonstrator’s performance in smooth, regular MDPs with sublinear $\gamma_{(Hn)}$ (for a formal definition, we refer to Lemma A.6 in the Appendix). We can also prove a more general result under weaker assumptions: as long as the policy is regular, the imitator will reach the noisy expert performance in arbitrary, non-smooth MDPs, albeit only in expectation. A full derivation of this further result can be found in Appendix B.
@@ 4.2 Practical Algorithms
Theorem 4.2 guarantees that, under regularity assumptions, the adaptive demonstration sampling scheme leads to convergence of the imitator’s performance to the optimal one. However, this criterion involves state occupancies and a conditional entropy term, which are hard to access or estimate in practice. Thus, here we derive a practical objective to be deployed in general settings. We first rephrase the objective from Equation 1 in its entropy form:
$c_{n}=$ $\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1}),$
with $c\sim\mu_{c},(s_{0},\dots)\sim\bm{\tau}(c),\tau^{\prime}\sim\bm{\tau}(c^{\prime}),$
$\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}).$ (2)
@@
Computing the expectation over a policy’s occupancy over states or trajectories is in general intractable in continuous state spaces. Fortunately, a coarse empirical estimate can be obtained as soon as few expert demonstrations become available. The expectation $\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}(\cdot)$ can be estimated through a single sample, which is always available in the form of the trajectories $\hat{\tau}_{1:n-1}$ collected so far, as they have effectively been sampled from $\bm{\tau}(c_{1:n-1})$. However, the remaining two expectations (i.e., $\mathop{\mathbb{E}}_{\tau^{\prime}\sim\bm{\tau}(c^{\prime})}(\cdot)$ and $\mathop{\mathbb{E}}_{(s_{0},\dots)\sim\bm{\tau}(c)}(\cdot)$) involve the distribution over trajectories for an arbitrary task, which might not have been demonstrated yet. We observe that, at round $n$, the tasks demonstrated so far induce the empirical distribution ${\hat{\mu}_{c}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{c_{i}}(\cdot)}$, while the trajectories collected similarly induce ${\bm{\hat{\tau}}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{\hat{\tau}_{i}}(\cdot)}$, where $\delta$ indicates the Dirac delta distribution. We can show that expectations over the trajectory distribution for an arbitrary task $c\in\mathcal{C}$ can be estimated through importance sampling (i.e., by sampling trajectories from $\bm{\hat{\tau}}(\cdot)$ instead of $\bm{\tau}(\cdot|c)$): ${\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(\cdot|c)}f(\tau)=\mathop{\mathbb{E}}_{\tau\sim\bm{\hat{\tau}}(\cdot)}\frac{\bm{\tau}(\tau|c)}{\bm{\hat{\tau}}(\tau)}f(\tau)}$. The importance weights can then be estimated as
@@
where $\tau=(s_{0},a_{0},\dots)$ and $\bm{\tilde{\pi}}$ can be approximated with the current estimate of $\bm{\pi}$. Intuitively, the likelihood ratio of a trajectory under two different tasks only depends on the likelihood of actions under the policy, and thus does not require knowledge of the MDP. As the estimate may be inaccurate for small numbers of samples, in practice the algorithm can invest the first few rounds to query a single demonstration for each of the tasks (in case they are countable and few) or to sample the task space uniformly. On the other hand, the high-variance of the estimate can be controlled by practical solutions such as clipping. We present the resulting empirical estimate for Equation 2 in full in Appendix C, and a qualitative analysis of importance weights in Appendix K.
@@
     if $n+1\;\%\;B=0$ then
        $\mathcal{D}_{n+1}=\mathcal{D}_{n+1-B}\cup\{c_{n-B+1:n},\tau_{n-B+1:n}\}$
        Update $\pi_{n+1}$ from $\pi_{n+1-B}$ with $\mathcal{D}_{n+1}$
     end if
@@
Thus, when the policy can be modeled as a GP, the only approximation needed concerns occupancy estimation. We refer to this first, practical instantiation as AMF-GP, and present a general algorithmic framework in Algorithm 1. Application of the method to policies parameterized by neural networks will adopt the same scheme. It will also require additional care on two distinct topics: kernel approximtions and mitigation of catastrophic forgetting.